40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Feb 7 21:47:44 2018, Craig, awade, DailyProgress, NoiseBudget, February 7th Noisebudget 20180207_221703noiseBudget.pdf
    Reply  Thu Feb 8 18:38:04 2018, rana, DailyProgress, NoiseBudget, February 7th Noisebudget 
       Reply  Fri Feb 9 11:36:08 2018, awade, DailyProgress, NoiseBudget, 200-3000 Hz hump search 
          Reply  Mon Feb 12 16:29:50 2018, awade, Craig, DailyProgress, NoiseBudget, 200-3000 Hz hump search 
             Reply  Fri Feb 16 12:41:32 2018, awade, Craig, DailyProgress, NoiseBudget, 200-3000 Hz hump search Beatnote_ASD_20180215_121922.pdfBeatnote_ASD_20180215_193422.pdfBeatnote_CurrentASD.pdf
Message ID: 2094     Entry time: Fri Feb 16 12:41:32 2018     In reply to: 2087
Author: awade, Craig 
Type: DailyProgress 
Category: NoiseBudget 
Subject: 200-3000 Hz hump search 

Following up on the last optimization of RFAM (see PSL:2087), I tried to obtain the -80 dBm RFAM at the north FSS refl PD with the flipper mirror up.  For some reason I am unable to reproduce the previous achieved RFAM lower bound (-80 dBm). 

We seem to be limited to -50 dBm with the flipper up.  I'm not sure what is going on here. Its possible that the large unfocused beam of the flipper reflection is givening secondary beams going through the Faraday isolator and we are actually looking at RF components mixed down with itself on a secondary reflection.  Also even after installing a PBS before the PMC there is still significant polarization drift.  The RFAM stays optimal for only 15 minutes or less. The EOM has an insulated hat, the waveplates and PMC do not. I don't know if we need to attempt a temperature control of these critical components.

Below I have attached a few ASDs taken with Criag's new live calibrated ASD generator (see PSL:2090).  One is taken before optimization of the RFAM on the north path, the next is imediatly after at 19:34 pm yesterday and the third is from this morning.  The system stopped logging BN spectra some time after 7 pm yesterday. When I checked early this morning at about 1 am it hump had returned.  I don't have the csv data files handy, so I am unable to complie them into a single plot, but you can see that the hump was greatly reduced with polarization optimization into the north 14.75 MHz EOM. 

It seems like this hump that pops up is a 14.75 MHz residual AM issue.  I'm guessing that the mechanism is that the lock point error offset is allowing intensity noise to couple into our measurement.  However, in the past Evan implemented an ISS only to find that it didn't improve the BN noise. A good test would be to inject an AM modulation line or swept sine and see how it propagates to the BN above what we would expect from photo thermal noise.   We need to address how to stabilize or remove this polarization drift. Active control sounds like a massive pain. Passive temperature stabilization might buy us an 30 to 60  minutes of good measurement but is not a good permanent solution.  Ideally we would locate the worst offending optic that is causing drift and deal with that.

Edit awade Fri Feb 16 13:07:12 2018: staring at the second attachment Beatnote_ASD_20180215_193422.pdf, something doesn't seem quite right. We definitely saw 0.3-0.6 Hz/sqrtHz on the SR785 last night. However, the roll up below 100 Hz in the auto generated ASD doesn't seem to have the same characteristic shape for scatter that we were looking at before.  Craig was working on scatter in the transmission table but the scatter shelf was back this morning so not likely that we really found the worst offending scatter.  Also, not sure if we are limited by the ADC noise above a 1 kHz there.  We need to go back an measure this with the SR785 and also check the PLL loop gain.  

Attachment 1: Beatnote_ASD_20180215_121922.pdf  213 kB  | Hide | Hide all
Beatnote_ASD_20180215_121922.pdf
Attachment 2: Beatnote_ASD_20180215_193422.pdf  208 kB  | Hide | Hide all
Beatnote_ASD_20180215_193422.pdf
Attachment 3: Beatnote_CurrentASD.pdf  211 kB  | Hide | Hide all
Beatnote_CurrentASD.pdf
ELOG V3.1.3-